# Strict Optimality of Frequency Estimation Under Local Differential Privacy

### 12 Mar 2026 v1

@@ 2.2 Frequency Estimation
[…]
Assuming the estimator is unbiased (i.e., $\forall x:\mathbb{E}[\hat{f}(x)]=f(x)$), we have $\mathbb{E}[(\hat{f}(x)-f(x))^{2}]=\text{Var}(\hat{f}(x))$. Consequently, the $\mathcal{L}_{2}$ loss can be expressed as:
$\mathcal{L}_{2}(\hat{f})=\max_{\mathbf{X}:|\mathbf{X}|=n}\sum_{x=1}^{d}\text{Var}(\hat{f}(x)).$ (1)
The relationship between $\mathcal{L}_{1}$ and $\text{Var}(\hat{f}(x))$ is established in the following theorem:
Theorem 2.3.
For a sufficiently large dataset size $n$, the $\mathcal{L}_{1}$ loss can be approximated as:
$\mathcal{L}_{1}(\hat{f})\approx\max_{\mathbf{X}:|\mathbf{X}|=n}\sum_{x\in[d]}\sqrt{\frac{2}{\pi}\text{Var}(\hat{f}(x))}.$ (2)
The proof is provided in Appendix A.2.
@@ 2.3 Extremal Configuration
Any frequency estimator lacking an extremal configuration can be transformed into one that possesses such a configuration.
@@
During frequency reconstruction, the estimator can simply replace any occurrence of $o_{1}$ or $o_{2}$ as the response $o$, rendering it identical to the original reconstruction. Consequently, an optimized frequency estimator with an extremal configuration always exists.
@@ 2.4 Symmetric Configuration
where $p^{*}$ and $q^{*}$ are constants independent of $x$ and $x^{\prime}$, and SS is the support matrix of $\mathcal{M}$. The values $p^{*}$ and $q^{*}$ are referred to as the self-support and cross-support probabilities, respectively.
@@
where $P$ is the perturbation matrix of $\mathcal{M}$, $I$ is the identity matrix, and $J$ is the all-ones matrix. $S^{T}P$ is referred to as the support probability matrix.
A reconstruction of frequency $f(x)$ is proposed by [wang2017localhashing] as:
$\hat{f}(x)=\frac{(\sum_{i\in[n]}S_{O^{(i)},x})/n-q^{*}}{p^{*}-q^{*}},$ (5)
where $S_{O^{(i)},x}$ is an indicator function denoting whether the response $O^{(i)}$ from the $i$-th object supports the value $x$. This estimation can also be represented in matrix form:
@@
$\mathbb{E}[\hat{f}(x)]=f(x)$ (7)
$\text{Var}(\hat{f}(x))=\frac{q^{*}(1-q^{*})}{n(p^{*}-q^{*})^{2}}+f(x)\frac{1-p^{*}-q^{*}}{n(p^{*}-q^{*})}.$ (8)
This theorem was originally proven in [wang2017localhashing]. The result can be verified by treating the term $\sum_{i\in[n]}S_{o^{(i)},x}$ as the sum of two independent binomial random variables: $\text{Bin}(nf(x),p^{*})$ and $\text{Bin}(n(1-f(x)),q^{*})$. These variables represent the number of responses from the original value $x$ and supporting $x$, and the number of responses from the original value other than $x$ and supporting $x$, respectively.
Further in this section, we provide a proof for the following claim:
Claim 2.1.
There exists a frequency estimator with a symmetric configuration that achieves optimality under both $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ losses.
The proof proceeds by transforming any frequency estimator that does not satisfy the symmetric configuration into one that does. We then demonstrate that the resulting estimator matches or exceeds the original’s performance in terms of worst-case precision.
Let a frequency estimator be denoted by $\hat{f}=F(P,Q)$, with its underlying perturbation mechanism denoted by $\mathcal{M}$. We construct a new perturbation mechanism $\mathcal{U}$ as follows:
@@
where $a$ and $b$ are constants, aligned with Definition 2.5. Meanwhile, an unbiased reconstruction matrix for $\mathcal{U}$ can be derived from the original reconstruction matrix $Q$:
@@
Proof: Given $P_{i}=PZ_{i}$, it follows that $Q_{i}P_{i}=I$. Since each permutation is sampled uniformly at random, the perturbation matrix for $\mathcal{U}$ is given by:
@@
As a result, a frequency estimator based on a uniformly random permutation is formally defined as follows:
@@
where $\hat{f}_{U}=\text{URP}(P,Q)$ and $\hat{f}_{i}=F(P_{i},Q_{i})$, respectively.
@@
The variance of $\hat{f}_{U}=\text{URP}(P,Q)$ satisfies
@@
Given that the permutation does not affect the worst-case scenario, the worst-case precision of $F(P,Q)$ also applies to $\text{URP}(P,Q)$, so Lemma 2.3 implies the following:
Corollary 2.1.
[…]
where $\hat{f}=F(P,Q)$ is and $\hat{f}_{U}=\text{URP}(P,Q)$.
If $\hat{f}$ is an optimal frequency estimator, then $\hat{f}_{U}$ must also be optimal, which concludes the proof of the following:
Theorem 2.9.
Claim 2.1 is correct.
@@
For a frequency estimator with a uniformly random permutation:
$\mathcal{L}_{1}(\hat{f})\approx\sqrt{\frac{2d}{\pi}\mathcal{L}_{2}(\hat{f})}$.
@@ 3 Strict Optimality in Precision
The proven Claim 2.1 suggest the existence of an optimal frequency estimator based on uniformly random perturbation (URP).
Corollary 2.2 demonstrates that optimizing the $\mathcal{L}_{2}$ loss for a URP frequency estimator simultaneously optimizes its $\mathcal{L}_{1}$ loss. Consequently, this section focuses on minimizing the $\mathcal{L}_{2}$ loss for a URP frequency estimator $\hat{f}_{U}=\text{URP}(P,Q)$. Following Theorem 2.8, we have:
$\mathcal{L}_{2}(\hat{f}_{U})=\sum_{x}(\alpha-\beta)f(x)+\beta-f(x)=\alpha+(d-1)\beta-1\\ =\frac{1}{d}\sum_{i}[(Q\circ Q)P]_{ii}+\frac{1}{d}\sum_{i}\sum_{j\neq i}[(Q\circ Q)P]_{ij}-1\\ =\frac{1}{d}\sum_{i}\sum_{j}[(Q\circ Q)P]_{ij}-1=\|Q\overline{P}^{\frac{1}{2}}\|_{2}-1.$
where $\overline{P}=\text{diag}(\frac{1}{d}P\mathbf{1})$, and $\mathbf{1}$ denotes an all-ones column vector of size $d$; $\|\cdot\|_{2}^{2}$ represents the squared $L_{2}$ norm (i.e., the squared Frobenius norm). Therefore, optimizing $\mathcal{L}_{2}$ is equivalent to minimizing $\|Q\overline{P}^{\frac{1}{2}}\|_{2}^{2}$. The methodology employed in this section to derive the strict lower bound of $\mathcal{L}_{2}$ is as follows:
  1. 1.
    Find an optimized reconstruction matrix $Q^{*}$ that minimizes $\|Q\overline{P}^{\frac{1}{2}}\|_{2}$ as a function of $P$.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
    Minimize the $\mathcal{L}_{2}$ loss of the OSC with respect to the support size $k$ to establish the strict lower bound for all frequency estimators.
$\forall P,Q:\mathcal{L}_{2}(F(P,Q))\geq\mathcal{L}_{2}(\text{URP}(P,Q))\\ \geq\mathcal{L}_{2}(\text{URP}(P,Q^{*}))\geq\mathcal{L}_{2}(\text{URP}(P^{*},Q^{*}))\\ =\mathcal{L}_{2}(\text{OSC}(d,\epsilon,n,k))\geq\min_{k}\mathcal{L}_{2}(\text{OSC}(d,\epsilon,n,k)).$
We begin with the first step:
@@
Let $\hat{f}_{U}=\text{URP}(P,Q)$. For a fixed perturbation matrix $P$, the $\mathcal{L}_{2}$ loss is minimized when
@@
$\min_{Q}\mathcal{L}_{2}(\hat{f}_{U})=\text{Tr}((P^{T}\overline{P}^{-1}P^{T})^{-1})-1,$
@@
$\min_{Q}\mathcal{L}_{2}\\ =\frac{1}{d}\sum_{i}\sum_{j}[(Q^{*}\circ Q^{*})P]_{ij}-1=\frac{1}{d}\sum_{i}[(Q^{*}\circ Q^{*})P\mathbf{1}]_{i}-1\\ =\text{Tr}(Q^{*}\overline{P}(Q^{*})^{T})-1=\text{Tr}((P^{T}\overline{P}^{-1}P)^{-1})-1.\qed$
With the optimized $Q^{*}$, we now proceed to find the optimal $P$ to minimize $\mathcal{L}_{2}$. Recall that Lemma 2.1 establishes that an extremal configuration is characterized by a constant self-support probability $p^{*}$, such that $\forall x:\sum_{o}e^{\epsilon}S_{o,x}p_{o}=p^{*}$. Building on this, we provide the following result:
@@
$\mathcal{L}_{2}=\frac{1}{d}+\sum_{i=2}^{d}\frac{1}{\lambda_{i}}\geq\frac{1}{d}+\frac{(d-1)^{2}}{\Lambda}\geq\frac{1}{d}+\frac{(d-1)^{2}}{\max\Lambda}.$
@@
Definiton 3.1.
A frequency estimator has an Optimal Symmetric Configuration (OSC) if it has an extremal and symmetric configuration with constant support, and employs Eq. (5) to reconstruct.
Theorems 3.2, 3.3, and 2.6 establish that a URP frequency estimator achieves an OSC. Thus, we have:
$\min_{\hat{f}_{\text{OSC}}}\mathcal{L}_{2}(\hat{f}_{\text{OSC}})\leq\min_{\hat{f}_{\text{URP}}}\mathcal{L}_{2}(\hat{f}_{\text{URP}}).$
Meanwhile, since $\min_{\hat{f}_{\text{URP}}}\mathcal{L}_{2}(\hat{f}_{\text{URP}})$ represents an optimized frequency estimator, it follows that $\min_{\hat{f}_{\text{OSC}}}\mathcal{L}_{2}(\hat{f}_{\text{OSC}})$ likewise corresponds to an optimized estimator.
To minimize $\mathcal{L}_{2}$ of an OSC estimator, we first formulate the variance of Eq. (5) as a function of the dictionary size $d$, privacy parameter $\epsilon$, dataset size $n$, and support size $k$. The extremal configuration ensures that $k$ can be directly derived from $p^{*}$, as shown in Lemma 2.1. Furthermore, the cross-support probability $q^{*}$ of OSC is derived as follows:
@@
$\mathcal{L}_{1}^{*}=\min_{k}\mathcal{L}_{1}(\hat{f})\approx\frac{\sqrt{2(d-1)[4de^{\epsilon}-(e^{\epsilon}+1)^{2}}]}{\pi\sqrt{n}(e^{\epsilon}-1)}.$ (12)
$\mathcal{L}_{2}^{*}=\min_{k}\mathcal{L}_{2}(\hat{f})=\frac{(d-1)[4de^{\epsilon}-(e^{\epsilon}+1)^{2}]}{nd(e^{\epsilon}-1)^{2}}$ (13)
The detailed proof is provided in Appendix B.4. The core methodology involves finding the minima through differentiation with respect to $k$. Although this optimality is derived within the URP framework, any frequency estimator can be optimized by satisfying these properties:
@@
An frequency estimator is $(\mathcal{L}_{1},\mathcal{L}_{2})$-optimal if it has an optimal symmetric configuration with $k=\frac{d}{e^{\epsilon}+1}$.
@@
$\min_{\hat{f}_{OSC}}\mathcal{L}_{1}(\hat{f}_{OSC})\leq\mathcal{L}_{1}^{*}\sqrt{1+\alpha},$
$\min_{\hat{f}_{OSC}}\mathcal{L}_{2}(\hat{f}_{OSC})\leq\mathcal{L}_{2}^{*}(1+\alpha),$
@@
$\min_{\hat{f}}\mathcal{L}_{1}(\hat{f})\approx\frac{\sqrt{2d(d-1)(d+2e^{\epsilon}-2)}}{\pi\sqrt{n}(e^{\epsilon}-1)}.$ (14)
$\min_{\hat{f}}\mathcal{L}_{2}(\hat{f})=\frac{(d-1)(d+2e^{\epsilon}-2)}{n(e^{\epsilon}-1)^{2}}$ (15)

### 22 Mar 2026 v2

@@ 2.3 Distribution Estimation
The detailed proof is provided in Appendix A.2.
@@ 2.4 Extremal Configuration
Any estimator lacking an extremal configuration can be transformed into one that possesses such a configuration.
@@
An estimator can simply replace any occurrence of $o_{1}$ or $o_{2}$ as the response $o$, rendering it identical to the original reconstruction. Consequently, an optimal estimator with an extremal configuration always exists.
@@ 2.5 Symmetric Configuration
where $p^{*}$ and $q^{*}$ are constants independent of $x$ and $x^{\prime}$, and SS is the support matrix of $M$. The values $p^{*}$ and $q^{*}$ are referred to as the self-support and cross-support probabilities, respectively.
@@
where $P$ is the perturbation matrix of $M$, II is the identity matrix, and $J$ is the all-ones matrix. $S^{T}P$ is referred to as the support probability matrix.
An estimator for symmetric configuration is proposed by [wang2017localhashing] as:
$\hat{f}(x)=\hat{\theta}(x)=\frac{(\sum_{i\in[n]}S_{O^{(i)},x})/n-q^{*}}{p^{*}-q^{*}},$ (5)
where $S_{O^{(i)},x}$ is an indicator function denoting whether the response $O^{(i)}$ from the $i$-th object supports the value $x$. The support counts can also be represented in matrix form:
@@
$\mathbb{E}[\hat{f}]=\mathbf{f}$ (7)
$\text{Var}(\hat{f})=\frac{q^{*}(1-q^{*})}{n(p^{*}-q^{*})^{2}}+\mathbf{f}\frac{1-p^{*}-q^{*}}{n(p^{*}-q^{*})}.$ (8)
$\text{Var}(\hat{\theta})=\frac{q^{*}(1-q^{*})}{n(p^{*}-q^{*})^{2}}+\bm{\theta}\frac{1-p^{*}-q^{*}}{n(p^{*}-q^{*})}+\frac{\bm{\theta}-\bm{\theta}^{2}}{n}.$ (9)
[…]
2.6 Uniformly Random Permutation
This section presents an approach to transform an arbitrary linear estimator into the one with a symmetric configuration, which will be proven to be an optimal estimator in the following sections.
Let a linear estimator be denoted by $\hat{f}=\hat{\theta}=\text{LN}(P,Q)$, with its underlying perturbation mechanism denoted by $M$. We construct a new perturbation mechanism $U$ as follows:
@@ 2.6 Uniformly Random Permutation
where $a$ and $b$ are constants, aligned with Definition 2.5. Meanwhile, an unbiased reconstruction matrix for $U$ can be derived from the original reconstruction matrix $Q$:
@@
Proof: Given $P_{i}=PZ_{i}$, it follows that $Q_{i}P_{i}=I$. Since each permutation is sampled uniformly at random, the perturbation matrix for $U$ is given by:
@@
As a result, an estimator based on a uniformly random permutation is formally defined as follows:
@@
$\text{Var}(\hat{\theta}_{U})=(Q_{U}\circ Q_{U})P_{U}\bm{\theta}-\bm{\theta}^{2}=\frac{1}{D}\sum_{i\in[D]}\text{Var}(\hat{\theta}_{i})$ (12)
where $\hat{f}_{U}=\text{URP}(P,Q)$ and $\hat{f}_{i}=\hat{\theta}_{i}=\text{LN}(P_{i},Q_{i})$, respectively.
@@
The variance of $\text{URP}(P,Q)$ satisfies
@@
$\text{Var}(\hat{\theta}_{U})$ $=(\alpha-\beta)\bm{\theta}+\beta-\bm{\theta}^{2},$
@@
For a frequency or distribution estimator $\hat{r}$ with a uniformly random permutation:
$\lim_{n\rightarrow\infty}\mathcal{L}_{1}(\hat{r},n)=\sqrt{\frac{2d}{\pi}\mathcal{L}_{2}(\hat{r},n)}.$
@@ 3 Strict Optimality in Precision
This section frames the optimality of $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ as two distinct minimization problems:
$\min_{\hat{r}}\mathcal{L}_{1|2}(\hat{r},n)=\min_{P}\min_{\hat{r}\in\mathcal{R}(P)}\mathcal{L}_{1|2}(\hat{r},n),$ (13)
where $\hat{r}$ denotes either $\hat{f}$ or $\hat{\theta}$, and $\mathcal{R}(P)$ represents the set of all estimators that use matrix $P$ for perturbation ($\mathcal{R}$ stands for reconstruction). Subsequently, we define:
Definiton 3.1.
An estimator $\hat{r}\in\mathcal{R}(P)$ is $P$-optimal if and only if $\forall P:\mathcal{L}_{1|2}(\hat{r},n)=\min_{\hat{r}\in\mathcal{R}(P)}\mathcal{L}_{1|2}(\hat{r},n)$.
Sections 3.1 and 3.2 prove that a URP estimator can achieve $P$-optimality, while Section 3.3 minimizes its $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ to establish optimal bounds.
3.1 $P$-Optimal Distribution Estimator
We begin by analyzing $P$-optimality for distribution estimation. Since $\mathcal{L}_{1|2}$ is the supremum of $L_{1|2}$, it is naturally lower-bounded by any $L_{1|2}$:
$\min_{\hat{\theta}\in\mathcal{R}(P)}\mathcal{L}_{1|2}(\hat{\theta},n)\geq\min_{\hat{\theta}\in\mathcal{R}(P)}L_{1|2}(\hat{\theta},\bm{\theta},n).$
Subsequently, we prove that
@@ 3.1 $P$-Optimal Distribution Estimator
$\min_{\hat{\theta}\in\mathcal{R}(P)}L_{2}(\hat{\theta},\bm{\theta},n)\geq\frac{1}{n}[\text{Tr}(\mathscr{I}_{C}^{+})]\overset{\mathrm{def}}{=}L_{2}^{\downarrow}(P,\bm{\theta},n).$
$\lim_{n\rightarrow\infty}$ $\min_{\hat{\theta}\in\mathcal{R}(P)}L_{1}(\hat{\theta},\bm{\theta},n)\geq\sum_{x\in[d]}\sqrt{\frac{2}{n\pi}(\mathscr{I}_{C}^{+})_{x,x}}\overset{\mathrm{def}}{=}L_{1}^{\downarrow}(P,\bm{\theta},n).$
where $\mathscr{I}_{C}^{+}=(P^{T}\text{diag}(P\bm{\theta})^{-1}P)^{-1}+\bm{\theta}\bm{\theta}^{T}$ and $L_{1|2}^{\downarrow}$ represents the theoretical lower bound of $L_{1|2}$.
The proof, provided in Appendix B.1, is inspired by the derivation of the Cramér–Rao bound using Fisher information.
Consequently, We will establish the $P$-optimality of $\hat{\theta}_{U}=\text{URP}(P,Q)$ by proving:
$\min_{Q}\mathcal{L}_{1|2}(\text{URP}(P,Q),n)=L_{1|2}^{\downarrow}(P,\frac{1}{d}\mathbf{1},n)$ (14)
where $\mathbf{1}$ denotes an all-ones column vector of size $d$. Note that the equality of $\mathcal{L}_{1}$ holds asymptotically as $n\to\infty$. Referring to Theorem 2.9, we derive the $L_{2}$ norm of a URP estimator as:
$L_{2}(\hat{\theta},\bm{\theta},n)=\sum_{x}(\alpha-\beta)\theta_{x}+\beta-\theta_{x}^{2}=\alpha+(d-1)\beta-\bm{\theta}^{T}\bm{\theta}\\ =\frac{1}{d}\sum_{i}[(Q\circ Q)P]_{ii}+\frac{1}{d}\sum_{i}\sum_{j\neq i}[(Q\circ Q)P]_{ij}-\bm{\theta}^{T}\bm{\theta}\\ =\frac{1}{d}\sum_{i}\sum_{j}[(Q\circ Q)P]_{ij}-\bm{\theta}^{T}\bm{\theta}=\|Q\overline{P}^{\frac{1}{2}}\|_{2}-\bm{\theta}^{T}\bm{\theta}.$ (15)
where $\overline{P}=\text{diag}(\frac{1}{d}P\mathbf{1})$; $\|\cdot\|_{2}^{2}$ represents the squared $L_{2}$ norm of a matrix (i.e., the squared Frobenius norm). It follows that:
$L_{2}(\hat{\theta}_{U},\bm{\theta},n)\leq L_{2}(\hat{\theta}_{U},\frac{1}{d}\mathbf{1},n)=\|Q\overline{P}^{\frac{1}{2}}\|_{2}-\frac{1}{d}=\mathcal{L}_{2}(\hat{\theta}_{U},n).$
Optimizing $\mathcal{L}_{2}$ is thus equivalent to minimizing $\|Q\overline{P}^{\frac{1}{2}}\|_{2}$. For the reconstruction matrix, we have:
Theorem 3.2.
Let $\hat{\theta}_{U}=\text{URP}(P,Q)$. For a fixed perturbation matrix $P$, the $\mathcal{L}_{2}$ loss is minimized when
@@
$\min_{Q}\mathcal{L}_{2}(\hat{\theta}_{U},n)=\frac{1}{n}[\text{Tr}((P^{T}\overline{P}^{-1}P^{T})^{-1})-\frac{1}{d}],$
@@
$\min_{Q}\mathcal{L}_{2}(\hat{\theta}_{U},n)\\ =\frac{1}{d}\sum_{i}\sum_{j}[(Q^{*}\circ Q^{*})P]_{ij}-\frac{1}{d}=\frac{1}{d}\sum_{i}[(Q^{*}\circ Q^{*})P\mathbf{1}]_{i}-\frac{1}{d}\\ =\text{Tr}(Q^{*}\overline{P}(Q^{*})^{T})-\frac{1}{d}=\text{Tr}((P^{T}\overline{P}^{-1}P)^{-1})-\frac{1}{d}.\qed$
This proves the $\mathcal{L}_{2}$ component of Eq. (14). Following Theorem 2.3, we obtain:
$\lim_{n\rightarrow\infty}\mathcal{L}_{1}(\text{URP}(P,Q^{*}),n)=L_{1}^{\downarrow}(P,\frac{1}{d}\mathbf{1},n).$ (16)
Consequently, we conclude:
Theorem 3.3.
$\text{URP}(P,Q^{*})$ is a $P$-optimal distribution estimator.
3.2 $P$-Optimal Frequency Estimator
We now extend the $P$-optimality established for distribution estimation to frequency estimation. First, consider the following sampling property:
Theorem 3.4.
Assume a dataset contains $n$ objects sampled i.i.d. from a categorical distribution $\mathcal{C}(\bm{\theta})$ and subsequently perturbed by an LDP mechanism. If the same unbiased estimator $\hat{f}=\hat{\theta}$ is employed for both frequency and distribution estimation, it satisfies:
$\sum_{\mathbf{f}}\Pr(\mathbf{f}|\bm{\theta})\text{Var}(\hat{f}|\mathbf{f})=\text{Var}(\hat{\theta})+\frac{\bm{\theta}^{2}-\bm{\theta}}{n}.$
The proof is provided in Appendix B.2. By relating these variances into $L_{2}$ and applying Theorem 3.1, we obtain:
$\sum_{\mathbf{f}}\Pr(\mathbf{f}|\bm{\theta})L_{2}(\hat{f},\mathbf{f},n)=L_{2}(\hat{\theta},\bm{\theta},n)+\frac{\bm{\theta}^{T}\bm{\theta}-1}{n}\\ \geq\frac{1}{n}[\text{Tr}((P^{T}\text{diag}(P\bm{\theta})^{-1}P)^{-1})-1].$
Regarding frequency estimation, Eq. (15) is adjusted to $\mathcal{L}_{2}(\hat{f}_{U})=\|Q\overline{P}^{\frac{1}{2}}\|_{2}-1$, which modifies Theorem 3.2 as follows:
$\mathcal{L}_{2}(\hat{f}_{U}=\text{URP}(P,Q^{*}),n)=\frac{1}{n}[\text{Tr}((P^{T}\overline{P}^{-1}P)^{-1})-1],$
where $\overline{P}=\text{diag}(P\frac{1}{d}\mathbf{1})$. Since the worst-case loss $\mathcal{L}_{2}$ is lower-bounded by the average $L_{2}$ loss, we have:
$\min_{\hat{f}\in\mathcal{R}(P)}\mathcal{L}_{2}(\hat{f},n)\geq\min_{\hat{f}\in\mathcal{R}(P)}\sum_{\mathbf{f}}\Pr(\mathbf{f}|\bm{\theta}=\frac{1}{d}\mathbf{1})L_{2}(\hat{f},\mathbf{f},n)\\ \geq\frac{1}{n}[\text{Tr}((P^{T}\overline{P}^{-1}P)^{-1})-1]=\mathcal{L}_{2}(\text{URP}(P,Q^{*}),n),$
By definition, $\mathcal{L}_{2}(\text{URP}(P,Q^{*}),n)\geq\min_{\hat{f}\in\mathcal{R}(P)}\mathcal{L}_{2}(\hat{f},n)$, implying that these quantities must be identical. Thus, $\text{URP}(P,Q^{*})$ is also $P$-optimal for frequency estimation. Similarly to Eq. (16), this $P$-optimality extends to $\mathcal{L}_{1}$. That is,
Theorem 3.5.
$\text{URP}(P,Q^{*})$ is a $P$-optimal frequency estimator.
3.3 Strict Lower Bound
Since $\text{URP}(P,Q^{*})$ is $P$-optimal for both frequency and distribution estimation, Eq. (13) ensures that $\min_{P}\mathcal{L}_{1|2}(\text{URP}(P,Q^{*}),n)$ constitutes the strict lower bound for $\mathcal{L}_{1|2}$. This section derives these values as follows:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Minimize the $\mathcal{L}_{2}$ loss of the OSC with respect to the support size $k$ to establish the strict lower bound for all estimators.
$\forall P:\mathcal{L}_{2}(\text{URP}(P,Q^{*}))\geq\mathcal{L}_{2}(\text{URP}(P^{*},Q^{*}))\\ =\mathcal{L}_{2}(\text{OSC}(d,\epsilon,n,k))\geq\min_{k}\mathcal{L}_{2}(\text{OSC}(d,\epsilon,n,k)).$
Notably, this section focuses on frequency estimation, but all results are applicable to distribution estimation by adjusting the last term of their variance expressions, i.e., by replacing $\mathbf{f}$ with $\bm{\theta}^{2}$. Consequently, $\mathcal{L}_{2}(\hat{\theta},n)=\mathcal{L}_{2}(\hat{f},n)+\frac{1-1/d}{n}$.
We now proceed to find the optimal $P$ to minimize $\mathcal{L}_{2}$. Recall that Lemma 2.1 establishes that an extremal configuration is characterized by a constant self-support probability $p^{*}$, such that $\forall x:\sum_{o}e^{\epsilon}S_{o,x}p_{o}=p^{*}$. Building on this, we provide the following result:
@@ 3.3 Strict Lower Bound
$\text{Tr}(P^{T}\overline{P}^{-1}P)=\frac{1}{d}+\sum_{i=2}^{d}\frac{1}{\lambda_{i}}\geq\frac{1}{d}+\frac{(d-1)^{2}}{\Lambda}\geq\frac{1}{d}+\frac{(d-1)^{2}}{\max\Lambda}.$
@@
Definiton 3.2.
An estimator has an Optimal Symmetric Configuration (OSC) if it has an extremal and symmetric configuration with constant support, and employs Eq. (5) to reconstruct the frequencies or distributions.
Theorems 3.6, 3.7, and 2.7 establish that a URP estimator achieves an OSC. Thus, we have:
$\min_{\hat{f}_{\text{OSC}}}\mathcal{L}_{2}(\hat{f}_{\text{OSC}},n)\leq\min_{\hat{f}_{\text{URP}}}\mathcal{L}_{2}(\hat{f}_{\text{URP}},n).$
Furthermore, since $\min_{\hat{f}_{\text{URP}}}\mathcal{L}_{2}(\hat{f}_{\text{URP}})$ has been established as an optimal bound, it follows that $\min_{\hat{f}_{\text{OSC}}}\mathcal{L}_{2}(\hat{f}_{\text{OSC}})$ similarly constitutes an optimal bound.
To derive $\mathcal{L}_{2}$ of an OSC estimator, we first formulate the variance of Eq. (5) as a function of the dictionary size $d$, privacy parameter $\epsilon$, dataset size $n$, and support size $k$. The extremal configuration ensures that $k$ can be directly derived from $p^{*}$, as shown in Lemma 2.1. Furthermore, the cross-support probability $q^{*}$ of OSC is derived as follows:
@@
$\mathcal{L}_{1}^{*}(n)=\min_{k}\mathcal{L}_{1}(\hat{f},n)\approx\frac{\sqrt{2(d-1)[4de^{\epsilon}-(e^{\epsilon}+1)^{2}}]}{\pi\sqrt{n}(e^{\epsilon}-1)}.$ (18)
$\mathcal{L}_{2}^{*}(n)=\min_{k}\mathcal{L}_{2}(\hat{f},n)=\frac{(d-1)[4de^{\epsilon}-(e^{\epsilon}+1)^{2}]}{nd(e^{\epsilon}-1)^{2}}$ (19)
The detailed proof is provided in Appendix B.6. The core methodology involves finding the minima through differentiation with respect to $k$. Consequently, the derivation throughout this section establishes a sufficient condition for an optimal estimator:
@@
An estimator is $(\mathcal{L}_{1},\mathcal{L}_{2})$-optimal if it has an optimal symmetric configuration with $k=\frac{d}{e^{\epsilon}+1}$.
@@
$\min_{\hat{f}_{OSC}}\mathcal{L}_{1}(\hat{f}_{OSC},n)\leq\mathcal{L}_{1}^{*}(n)\sqrt{1+\alpha},$
$\min_{\hat{f}_{OSC}}\mathcal{L}_{2}(\hat{f}_{OSC},n)\leq\mathcal{L}_{2}^{*}(n)(1+\alpha),$
@@
$\min_{\hat{f}}\mathcal{L}_{1}(\hat{f},n)\approx\frac{\sqrt{2d(d-1)(d+2e^{\epsilon}-2)}}{\pi\sqrt{n}(e^{\epsilon}-1)}.$ (20)
$\min_{\hat{f}}\mathcal{L}_{2}(\hat{f},n)=\frac{(d-1)(d+2e^{\epsilon}-2)}{n(e^{\epsilon}-1)^{2}}$ (21)